Data for Hermann Corporation are shown below:

Per Unit Percent of Sales
Selling price $ 70 100 %
Variable expenses 49 70
Contribution margin $ 21 30 %

Fixed expenses are $74,000 per month and the company is selling 4,400 units per month.

Required:

1-a. How much will net operating income increase (decrease) per month if the monthly advertising budget increases by $9,800 and monthly sales increase by $24,000?

1-b. Should the advertising budget be increased?
Answer #1
Current Proposed
Sales (70*4400)=308,000 (308,000+24,000)=332,000
Variable expenses (49*4400)=215600 (332,000*70%)=$232400
Contribution margin 92400 99600
Fixed expenses 74000 (74000+9800)=83800
Net operating income 18400 15800

Hence decrease in Net operating income=(18400-15800)=$2600

Hence advertising budget should not be increased.